OpenStudy (anonymous):

Write a possible explicit rule for the nth term of the sequence: 25, 15, 5, -5, -15,

OpenStudy (anonymous):

a = 25 d = 15 - 25 = 5-15 = -10 \[a_n = a+(n-1) d\]